OpenStudy (anonymous):

(Anyone pls help?) Quadrilateral LMNO has diagonals that intersect at point P. If LP=NP, MP = y + 25, and OP = 5y + 29, find the length of MO such that LMNO is a parallelogram. A.24 B.48 C.61 D.122

OpenStudy (anonymous):



OpenStudy (anonymous):

For a parallelogram we have \(MP=OP\) i.e. \(y+25=5y+29\). Can you solve for \(y\) (it's trivial)? Once you do, find \(MO=MP+OP=2(y+25)=\dots\)

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

the answers 24. Hey thanks!

OpenStudy (anonymous):

\(y=-1\) so we get \(MO=2(24)=48\)!

OpenStudy (anonymous):

\(24\) is just the length of *half* of \(MO\)!

OpenStudy (anonymous):